According to the U.S. Census Bureau, the two most popular female names in the United States are Mary and Patricia. Combined, these two names account for 3.7% of all female first names. The difference between the percentage of women named Mary and the percentage of women named Patricia are 1.5%. What percentage of all female first anmes in the United States are Mary and what percentage are Patricia?
Answer by stanbon(48567)

According to the U.S. Census Bureau, the two most popular female names in the United States are Mary and Patricia. Combined, these two names account for 3.7% of all female first names. The difference between the percentage of women named Mary and the percentage of women named Patricia are 1.5%. What percentage of all female first names in the United States are Mary and what percentage are Patricia?
----------------------------
Let M be % of Marys ;Let P be % of Patricias
---
Equation:
M+P = 0.037
M-P = 0.015
-------------------
Add to solve for "M"
2M = 0.052
M = 0.026 or 2.6% (% of Marys in the population)
---
Since M+P = 0.037, 0.026+P = 0.037
P = 1.1% (% of Patricias in the population)
================================================
Cheers,
Stan H.